# Nash equilibrium seeking for a class of quadratic-bilinear Wasserstein distributionally robust games

### v1

@@ Abstract
We consider a class of Wasserstein distributionally robust Nash equilibrium problems, where agents construct heterogeneous data-driven Wasserstein ambiguity sets using private samples and radii, in line with their individual risk-averse behaviour. By leveraging relevant properties of this class of games, we show that equilibria of the original seemingly infinite-dimensional problem can be obtained as a solution to a finite-dimensional Nash equilibrium problem. We then reformulate the problem as a finite-dimensional variational inequality and establish the connection between the corresponding solution sets. Our reformulation has scalable behaviour with respect to the data size and maintains a fixed number of constraints, independently of the number of samples. To compute a solution, we leverage two algorithms, based on the golden ratio algorithm. The efficiency of both algorithmic schemes is corroborated through extensive simulation studies on an illustrative example and a stochastic portfolio allocation game, where behavioural coupling among investors is modeled.
@@
Data-driven Nash equilibrium seeking Wasserstein ambiguity sets Heterogeneous uncertainty
@@ 1 Introduction
A wide range of applications, from smart grids Saad1 and communication networks Scutari to social networks Acemoglu2013 can be modelled as a collection of self-interested interacting decision makers optimizing different cost functions under operational constraints. Game theory Basar1 provides the fundamental theoretical framework for analyzing such systems. Although investigating deterministic games can be adequate in some case studies Scutari , Paccagnan2017 , most real-world applications involve decision making under uncertainty, which stresses the need for the inclusion of stochasticity in the existing models. Several studies have explored uncertainty within a game theoretic context, based on particular assumptions on the probability distribution Kouvaritakis , Singh and/or the properties of the uncertainty sample space Aghassi2006 ; FukuSOCCP .
@@
    We study a class of heterogeneous data-driven Wasserstein distributionally robust games, where each agent’s ambiguity set is centered around an empirical probability distribution based on their individual data, while the Wasserstein radius is also set by each individual agent. We reformulate the original game as a robust Nash equilibrium problem and establish the connection between the distributionally robust and robust Nash equilibria of the corresponding problems. For this class of games, we demonstrate that the inner maximization can be solved without the use of epigraphic variables kuhn2019wasserstein , pantazis2023_DRG , which in a game-theoretic setting can lead to unshared coupling constraints. As such, our approach decreases computational complexity significantly. To the best of our knowledge, this is the first distributionally robust game-theoretic reformulation that leads to data-scalable results by leveraging the structure of the problem at hand.
@@ 2 Problem formulation
$d_{W}(\mathbb{Q}_{1},\mathbb{Q}_{2}):=$ $\inf_{\Pi\in\pazocal{J}(\xi_{1}\sim\mathbb{Q}_{1},\xi_{2}\sim\mathbb{Q}_{2})}\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\xi_{1}-\xi_{2}\|\Pi(d\xi_{1},d\xi_{2}),$
@@ 2.1 Problem formulation
Consider a population of agents with index set $\mathscr{N}=\{1,\dots,N\}$. Each agent $i\in\mathscr{N}$, given the decisions of the other agents $x_{-i}$, solves the following optimization problem:
$\min_{x_{i}\in X_{i}}\max_{\mathbb{Q}_{i}\in\mathscr{P}_{i}}\{f_{i}(x_{i},x_{-i})+\mathbb{E}_{\xi_{i}\sim\mathbb{Q}_{i}}[g_{i}(x_{i},x_{-i},\xi_{i})]\},$
where $f_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n(N-1)}\rightarrow\mathbb{R}$, $g_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n(N-1)}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ for all $i\in\mathscr{N}$ and $\mathscr{P}_{i}$ is the ambiguity set of the uncertain parameter $\xi_{i}$. We call the collection of the coupled optimization problems above for all agents $i\in\mathscr{N}$ as game $G$. For game $G$, we define the notion of distributionally robust Nash equilibrium as follows:
@@
$\mathscr{P}_{i}=\{Q_{i}\in\mathscr{P}(\mathbb{R}^{m}):d_{W}(Q_{i},\hat{\mathbb{P}}_{K_{i}})\leq\varepsilon_{i}\},$ (2)
where $\mathscr{P}(\mathbb{R}^{m})$ denotes the collection of all probability distributions defined on the support set $\mathbb{R}^{m}$. We impose the following assumption:
@@
Let Assumption 1 hold. Fix the Wasserstein radii $\varepsilon_{i}$ and consider a multi-sample $\xi_{K_{i}}\in\mathbb{R}^{mK_{i}}$ for each agent $i\in\pazocal{N}$. Then, each optimization problem admits the following dual reformulation:
@@
$(x^{\ast}_{i},\lambda_{i}^{\ast})\in\operatorname*{arg\,min}_{x_{i}\in X_{i},\lambda_{i}\geq 0}J_{i}(x_{i},\lambda_{i},x^{\ast}_{-i})$
for all $i\in\mathscr{N}$, with $J_{i}$ as in (4). $\square$
@@ 2.2 Reformulation as a data-driven variational inequality problem
Proof: The proof follows from direct application of Proposition 1 by taking the pseudogradient of game $G_{R}$ considering that $\tilde{Q}_{i}$ is a diagonal matrix, hence $\frac{d\tilde{Q}_{i}(\lambda_{i})}{d\lambda_{i}}$ is obtained by differentiating the corresponding diagonal elements. $\blacksquare$
The resulting VI mapping can in general be nonmonotone. However, for a fixed set of best-response strategies $x^{\ast}_{-i}$, the resulting optimization problem for each agent $i\in\pazocal{N}$ is convex. This is an immediate result of the quadratic over linear structure of each optimization problem, whose Jacobian is positive semidefinite, given $x^{\ast}_{-i}$. Nonmonotonicity of the corresponding VI mapping implies that depending on the initialization point within the region $X$, different sets of equilibrium solutions may be reached with an equilibrium seeking algorithm. Note that those points satisfy the equilibrium condition of $G_{R}$.

### v2

@@ Abstract
We consider a class of Wasserstein distributionally robust Nash equilibrium problems, where agents construct heterogeneous data-driven Wasserstein ambiguity sets using private samples and radii, in line with their individual risk-averse preference. By leveraging relevant properties of this class of games, we show that equilibria of the original seemingly infinite-dimensional problem can be obtained as a solution to a finite-dimensional Nash equilibrium problem. We then reformulate the associated variational problem as a finite-dimensional variational inequality and establish the connection between the corresponding solution sets. Our reformulation has scalable behaviour with respect to the data size and maintains a fixed number of constraints, independently of the number of samples. To numerically compute a solution, we leverage two algorithms, based on the golden ratio algorithm. The efficiency of both algorithmic schemes is corroborated through extensive simulation studies on an illustrative example and a stochastic portfolio allocation game, where behavioural coupling among investors is modeled.
@@
Data-driven Nash Equilibrium Seeking Distributionally Robust Games Wasserstein Ambiguity sets Heterogeneous Uncertainty
@@ 1 Introduction
A wide range of applications, from smart grids Saad1 and communication networks Scutari to social networks Acemoglu2013 can be modeled as a collection of self-interested interacting decision makers optimizing different objective functions under operational constraints. Game theory Basar1 provides the fundamental theoretical framework for analyzing such systems. Although investigating deterministic games can be adequate in some case studies Scutari , Paccagnan2017 , most real-world applications involve decision making under uncertainty, which stresses the need for the inclusion of stochasticity in the existing models. Several studies have explored uncertainty within a game theoretic context, based on particular technical assumptions on the probability distribution Kouvaritakis , Singh and/or the properties of the uncertainty sample space Aghassi2006 ; FukuSOCCP .
@@
    We study a class of heterogeneous data-driven Wasserstein distributionally robust games with a quadratic-bilinear structure for each agent objective function. Each agent is considered to act as a selfish entity making individual decisions to optimize their own cost function, which depends on a private uncertain parameter. The game-theoretic formulation generalizes the optimization framework to allow for interdependence of the agent cost functions. Furthermore, as we model possibly heterogeneous probability distributions, we lift the assumption of a common ambiguity set, allowing each agent to construct their own, based on their private data and personalized Wasserstein radius. In this generalized setting, the solution set can significantly differ from those associated with common ambiguity sets mohajerin_esfahani_data-driven_2018 , netessine_wasserstein_2019 , shafiee2024 , Lanzetti2025
@@
    We reformulate the original game as a robust Nash equilibrium problem and establish the connection between the distributionally robust and robust Nash equilibria of the corresponding problems. For this class of games, we demonstrate that the inner maximization can be solved without the use of epigraphic variables kuhn2019wasserstein , pantazis2023_DRG , which, unlike distributionally robust optimization, in distributionally robust games can lead to unshared coupling constraints. As such, our approach decreases computational complexity significantly. To the best of our knowledge, this is the first distributionally robust game-theoretic reformulation that leads to data-scalable results by leveraging the structure of the problem at hand.
  3. (iii)
@@ 2 Problem Formulation
$d_{W}(\mathbb{Q}_{1},\mathbb{Q}_{2}):=$ $\left(\inf_{\Pi\in\pazocal{J}(\xi_{1}\sim\mathbb{Q}_{1},\xi_{2}\sim\mathbb{Q}_{2})}\int_{\mathbb{R}^{m}\times\mathbb{R}^{m}}\|\xi_{1}-\xi_{2}\|^{p}\Pi(d\xi_{1},d\xi_{2})\right)^{1/p},$
@@ 2.1 Problem Formulation
Consider a population of agents with index set $\mathscr{N}=\{1,\dots,N\}$. Based on that we define the following game:
$\forall i\in\mathscr{N}:\min_{x_{i}\in X_{i}}\max_{\mathbb{Q}_{i}\in\mathscr{P}_{i}}\{f_{i}(x_{i},x_{-i})+\mathbb{E}_{\xi_{i}\sim\mathbb{Q}_{i}}[g_{i}(x_{i},x_{-i},\xi_{i})]\},$
where $f_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n(N-1)}\rightarrow\mathbb{R}$, $g_{i}:\mathbb{R}^{n}\times\mathbb{R}^{n(N-1)}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ for all $i\in\mathscr{N}$ and $\mathscr{P}_{i}$ is the ambiguity set of the uncertain parameter $\xi_{i}$. We call the collection of the coupled optimization problems above for all agents $i\in\mathscr{N}$ as game ($G$).
For game $G$, we define the notion of distributionally robust Nash equilibrium as follows:
@@
$\mathscr{P}_{i}=\{Q_{i}\in\mathscr{P}(\mathbb{R}^{m})\mid d_{W}(Q_{i},\hat{\mathbb{P}}_{K_{i}})\leq\varepsilon_{i}\},$ (2)
where $\mathscr{P}(\mathbb{R}^{m})$ denotes the collection of all probability distributions defined on the support set $\mathbb{R}^{m}$.
2.2 Effect of heterogeneous ambiguity sets on the equilibrium set
We consider a simple example to illustrate how heterogeneity of each agents’ ambiguity sets, which correspond to an uncertain parameter $\xi\in\mathbb{R}^{2}$, can affect the solution set of DRNE even in the case of 1-Wasserstein distance $(p=1)$. In the following distributionally robust game, two self-interested agents solve the following interdependent optimization problems:
$\begin{cases}\min\limits_{x_{1}}\max\limits_{\mathbb{Q}_{1}\in\mathscr{P}_{\varepsilon_{1}}(\hat{\mathbb{P}}_{K_{1}})}\mathbb{E}_{\mathbb{Q}_{1}}[c_{11}x^{2}_{1}+c_{12}x_{1}x_{2}+\xi^{\top}x]\\ \min\limits_{x_{2}}\max\limits_{\mathbb{Q}_{2}\in\mathscr{P}_{\varepsilon_{2}}(\hat{\mathbb{P}}_{K_{2}})}\mathbb{E}_{\mathbb{Q}_{2}}[c_{21}x_{1}x_{2}+c_{22}x^{2}_{2}+\xi^{\top}x]\end{cases}$ (3)
The game in (3) can be equivalently written as follows netessine_wasserstein_2019 :
$\begin{cases}\min\limits_{x_{1}}c_{11}x_{1}^{2}+c_{12}x_{1}x_{2}+\frac{1}{K_{1}}\sum\limits_{k=1}^{K_{1}}(\xi^{(k)}_{1})^{\top}x+\varepsilon_{1}\|x\|_{*}\\ \min\limits_{x_{2}}c_{21}x_{1}x_{2}+c_{22}x_{2}^{2}+\frac{1}{K_{2}}\sum\limits_{k=1}^{K_{2}}(\xi^{(k)}_{2})^{\top}x+\varepsilon_{2}\|x\|_{*},\end{cases}$
[…]
$\begin{cases}2c_{11}\bar{x}_{1}+c_{12}\bar{x}_{2}+\varepsilon_{1}\partial_{x_{1}}\|\bar{x}\|_{*}=-\frac{1}{K_{1}}\sum\limits_{k=1}^{K_{1}}\xi^{(k)}_{1}\\ c_{21}\bar{x}_{1}+2c_{22}\bar{x}_{2}+\varepsilon_{2}\partial_{x_{2}}\|\bar{x}\|_{*}=-\frac{1}{K_{2}}\sum\limits_{k=1}^{K_{2}}\xi^{(k)}_{2}\end{cases}$ (4)
Considering the 2-norm for the transport cost, we have that:
Figure 1: Equilibrium solution for homogeneous (blue lines) and heterogeneous (red lines) ambiguity sets for a 2-player game with scalar decisions. The dashed lines correspond to the second equation in (5).
$\begin{cases}2c_{11}\bar{x}_{1}+c_{12}\bar{x}_{2}+\varepsilon_{1}\frac{\bar{x}_{1}}{\sqrt{\bar{x}_{1}^{2}+\bar{x}_{2}^{2}}}=-\frac{1}{K_{1}}\sum\limits_{k=1}^{K_{1}}\xi^{(k)}_{1}\\ c_{21}\bar{x}_{1}+2c_{22}\bar{x}_{2}+\varepsilon_{2}\frac{\bar{x}_{2}}{\sqrt{\bar{x}_{1}^{2}+\bar{x}_{2}^{2}}}=-\frac{1}{K_{2}}\sum\limits_{k=1}^{K_{2}}\xi^{(k)}_{2}\end{cases}$ (5)
We now show how, even for simple discrete distributions concentrated at one point, a difference between $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ can lead to a significant change to the equilibrium solution. Consider a discrete probability distribution $\mathbb{P}_{1}$ with full probability mass at $x=p_{1}$ (i.e., with probability 1) and 0 elsewhere, and the probability distribution $\mathbb{P}_{2}$ with full probability mass at $x=p_{2}$, 0 elsewhere:
$\begin{cases}2c_{11}\bar{x}_{1}+c_{12}\bar{x}_{2}+\varepsilon_{1}\frac{\bar{x}_{1}}{\sqrt{\bar{x}_{1}^{2}+\bar{x}_{2}^{2}}}=-p_{1}\\ c_{21}\bar{x}_{1}+2c_{22}\bar{x}_{2}+\varepsilon_{2}\frac{\bar{x}_{2}}{\sqrt{\bar{x}_{1}^{2}+\bar{x}_{2}^{2}}}=-p_{2}\end{cases}$ (6)
In Figure 1, we first consider $c_{11}=1$, $c_{12}=0.5$, $c_{21}=0.5$, $c_{22}=1$ and homogeneous ambiguity sets, i.e, $\varepsilon_{1}=0.1$, $\varepsilon_{2}=0.1$, $p_{1}=1$, $p_{2}=1$. Considering heterogeneous ambiguity sets ($\varepsilon_{1}=2$, $\varepsilon_{2}=0.1$, $p_{1}=1$, $p_{2}=-4$), we note that the equilibrium point changes. Thus, our setting generalizes standard distributionally robust optimization to games of agents with self-interested objectives that can be affected by different distributions, have their own individual data and select their own preferred risk-aversion via their Wasserstein radius. This example shows how different radii and different samples, can introduce information asymmetry in the problem at hand, thus significantly changing the set of equilibria.
3 Wasserstein Distributionally robust quadratic-bilinear games
In this work, we consider a class of quadratic-bilinear games under the 2-Wasserstein distance ($p=2$). We impose the following assumption:
@@ 3 Wasserstein Distributionally robust quadratic-bilinear games
Let Assumption 1 hold. Fix the Wasserstein radii $\varepsilon_{i}$ and consider a multi-sample $\xi_{K_{i}}\in\mathbb{R}^{mK_{i}}$ for each agent $i\in\pazocal{N}$. Then, each agent’s $i\in\pazocal{N}$ problem admits the following dual reformulation:
@@
$\forall i\in\pazocal{N}:(x^{\ast}_{i},\lambda_{i}^{\ast})\in\operatorname*{arg\,min}_{x_{i}\in X_{i},\lambda_{i}\geq 0}J_{i}(x_{i},\lambda_{i},x^{\ast}_{-i}),$
with $J_{i}$ as in (8). $\square$
@@ 3.1 Reformulation as a Data-driven Variational Inequality Problem
Proof: The proof follows by adaptation of Proposition 1 by taking the pseudogradient of game $G_{R}$ considering that $\tilde{Q}_{i}$ is a diagonal matrix, hence $\frac{d\tilde{Q}_{i}(\lambda_{i})}{d\lambda_{i}}$ is obtained by differentiating the corresponding diagonal elements. $\blacksquare$
3.2 Challenges of distributionally robust heterogeneous games compared to DRO
Let us illustrate the challenges that our game-theoretic problem entails compared to the standard DRO. Let us first define the associated distributionally robust optimization problem as follows:
$\min_{x\in X}\max_{\mathbb{Q}\in\mathscr{P}_{\epsilon}(\hat{\mathbb{P}}_{K})}\mathbb{E}_{\xi\sim\mathbb{Q}}[x^{\top}Qx+\xi^{\top}Q\xi+P(x)^{\top}\xi],$ (16)
First, note that the optimization problem in (16) is only a special case of the original game considered in (1). In particular, the setting in (1) considers more than one players, who make self-interested decisions based on their own private samples and different Wasserstein radii. The interdependence of each agent’s cost to the decisions of the other agents significantly affects the resulting decision, unlike optimization where this decision is collective and only nature is viewed as an adversarial player. This implies that the pseudogradient mapping might be non-monotone; On the contrary, in a standard optimization problem, the gradient map is always monotone.
For example, consider the following distributionally robust game with 2 agents $i\in\{1,2\}$ taking scalar decisions. Specifically, agent 1 aims at solving:
$\min_{x_{1}\in X_{1}}\max_{\mathbb{Q}_{1}\in\mathscr{P}_{1}}\mathbb{E}_{\xi_{1}\sim\mathbb{Q}_{1}}[q_{11}x^{2}_{1}+q_{12}x_{1}x_{2}+Q_{1}\xi^{2}_{1}+(a_{1}x_{1}+a_{2}x_{2})\xi_{1}],$
while agent 2 aims at solving:
$\min_{x_{2}\in X_{2}}\max_{\mathbb{Q}_{2}\in\mathscr{P}_{2}}\mathbb{E}_{\xi_{2}\sim\mathbb{Q}_{2}}[q_{22}x^{2}_{2}+q_{21}x_{1}x_{2}+Q_{2}\xi^{2}_{2}+(b_{1}x_{1}+b_{2}x_{2})\xi_{2}],$
By Theorem 1, the problem above can be written as:
$\forall\ i\in\mathscr{N}:\min_{(x_{i},\lambda_{i})\in X_{i}\times\mathbb{R}^{m}_{\geq 0}}J_{i}(x_{i},x_{j},\lambda_{i}),$ (17)
where
$J_{1}(x_{1},x_{2},\lambda_{1})=q_{11}x^{2}_{1}+q_{12}x_{1}x_{2}+\lambda_{1}\left(\epsilon^{2}_{1}-\frac{1}{K_{1}}\sum_{k_{1}=1}^{K_{1}}(\xi^{(k_{1})}_{1})^{2}\right)+\frac{1}{4K_{1}}\sum_{k_{1}=1}^{K_{1}}\frac{(a_{1}x_{1}+a_{2}x_{2}+2\lambda_{1}\xi^{(k_{1})}_{1})^{2}}{\lambda_{1}-\lambda_{max,1}},$
and
$J_{2}(x_{1},x_{2},\lambda_{2})=q_{22}x^{2}_{2}+q_{21}x_{1}x_{2}+\lambda_{2}\left(\epsilon^{2}_{2}-\frac{1}{K_{2}}\sum_{k_{2}=1}^{K_{2}}(\xi^{(k_{2})}_{2})^{2}\right)+\frac{1}{4K_{2}}\sum_{k_{2}=1}^{K_{2}}\frac{(b_{1}x_{1}+b_{2}x_{2}+2\lambda_{2}\xi^{(k_{2})}_{2})^{2}}{\lambda_{2}-\lambda_{max,2}}.$
To study the monotonicity of the VI mapping associated with (17), we can directly study whether the game Jacobian is positive semidefinite. The following lemma shows that for this class of games, even if $K_{1}=K_{2}$ (note that only the number is the same, thus still allowing for different samples per agent), the Jacobian is more complicated than the calculated one for standard optimization.
Lemma 4
Consider the functions
$G^{(k_{1})}_{1}(x_{1},x_{2},\lambda_{1})=2\xi^{(k_{1})}_{1}\lambda_{1,\max}+a_{1}x_{1}+a_{2}x_{2},$
$G^{(k_{2})}_{2}(x_{1},x_{2},\lambda_{2})=2\xi^{(k_{2})}_{2}\lambda_{2,\max}+b_{1}x_{1}+b_{2}x_{2},$
$\Delta_{1}=\lambda_{1}-\lambda_{1,\max}$
$\Delta_{2}=\lambda_{2}-\lambda_{2,\max},$ (18)
corresponding to each sample $k_{1}\in\{1,\dots,K_{1}\}$ and $k_{2}\in\{1,\dots,K_{2}\}$. Furthermore, consider for simplicity that $K_{1}=K_{2}=K$ and $q_{11}=q_{12}=q_{22}=q_{21}=0$. The symmetrized Jacobian $J$ is then given by
$J=\frac{1}{4K}\sum_{k=1}^{K}J^{(k)}_{s},$
where
$J^{(k)}_{s}=\begin{bmatrix}\frac{2a_{1}^{2}}{\Delta_{1}}&\frac{2a_{1}G^{(k)}_{1}(x,\lambda_{1})}{\Delta_{1}^{2}}&\frac{a_{1}a_{2}}{\Delta_{1}}+\frac{b_{1}b_{2}}{\Delta_{2}}&\frac{b_{1}G^{(k)}_{2}(x,\lambda_{2})}{\Delta_{2}^{2}}\\ \frac{2a_{1}G^{(k)}_{1}(x,\lambda_{1})}{\Delta_{1}^{2}}&\frac{2G^{(k)}_{1}(x,\lambda_{1})^{2}}{\Delta_{1}^{3}}&\frac{a_{2}G^{(k)}_{1}(x,\lambda_{1})}{\Delta_{1}^{2}}&0\\ \frac{a_{1}a_{2}}{\Delta_{1}}+\frac{b_{1}b_{2}}{\Delta_{2}}&\frac{a_{2}G^{(k)}_{1}(x,\lambda_{1})}{\Delta_{1}^{2}}&\frac{2b_{2}^{2}}{\Delta_{2}}&\frac{2b_{2}G^{(k)}_{2}(x,\lambda_{2})}{\Delta_{2}^{2}}\\ \frac{b_{1}G^{(k)}_{2}(x,\lambda_{2})}{\Delta_{2}^{2}}&0&\frac{2b_{2}G^{(k)}_{2}(x,\lambda_{2})}{\Delta_{2}^{2}}&\frac{2G^{(k)}_{2}(x,\lambda_{2})^{2}}{\Delta_{2}^{3}}\end{bmatrix}.$ (19)
Proof: The proof is obtained by performing standard algebraic operations on the defined Jacobian of the corresponding VI, considering the functions $G^{(k_{1})}$ and $G^{(k_{2})}$. We then use the assumption $K_{1}=K_{2}=K$ to pull a common sum outside and group the samples of the two agents in exhaustive pairs. $\blacksquare$
Based on Lemma 4, an important observation can then be made: When we consider only one agent, we derive as a special case of our original problem a distributionally robust optimization program with Jacobian:
$J^{(k)}_{s}=\begin{bmatrix}\frac{2a_{1}^{2}}{\Delta_{1}}&\frac{2a_{1}G^{(k_{1})}_{1}(x,\lambda_{1})}{\Delta_{1}^{2}}\\ \frac{2a_{1}G^{(k_{1})}_{1}(x,\lambda_{1})}{\Delta_{1}^{2}}&\frac{2G^{(k_{1})}_{1}(x,\lambda_{1})^{2}}{\Delta_{1}^{3}}\end{bmatrix},$ (20)
For any $k$, $J_{s}^{(k)}$ can then be shown to be positive semi-definite. Thus $J$ is positive semi-definite as the summation of positive semi-definite matrices. In contrast, the Jacobian (19) for the game in (16) might not be positive semi-definite which implies non-monotonicity. For example, the game in (18) with fixed decisions $x_{1}=2$, $x_{2}=2$, $\lambda_{1}=\lambda_{1,\max}+2$ and $\lambda_{2}=\lambda_{2,\max}+2$, and parameters $\lambda_{1,\max}=0.5$, $\lambda_{2,\max}=0.5$, $a_{1}=1$, $a_{2}=0.5$, $b_{1}=1$, $b_{2}=0.5$, has Jacobian with at least one negative eigenvalue meaning that the corresponding pseudogradient is not monotone.